# CausalMed: Causality-Based Personalized Medication Recommendation Centered on Patient health state

### v2

@@ Abstract.
submissionid: 137ccs: Information systems Information retrievalccs: Computing methodologies Causal reasoning and diagnosticsccs: Applied computing Health informatics
@@ 2.1. Medication Recommendation
Early works in medication recommendation like LEAP (Zhang et al., 2017) are instance-based, prescribing on single-visit data, and viewing medication recommendation as a multi-instance and multi-label classification task.
The first category of research focuses on modeling the sequential relationships in patients’ longitudinal medical records to address issues of personalization. RETAIN (Choi et al., 2016b) models patients’ histories to personalize disease predictions. MICRON (Yang et al., 2021a) adjusts medications for new symptoms to optimize therapy and minimize side effects. COGNet (Wu et al., 2022) translates patient’s diseases/procedures to medications and reuses effective past treatments.
The second category of methods aims to capture the relationships between entities, thereby addressing the problem of low accuracy in prescriptions. DMNC (Le et al., 2018) combines dynamic memory and deep learning to analyze case data and recommend precise medications. Several studies (Zheng et al., 2021b, 2023; Wu et al., 2023) utilize graph networks to describe the connections between medical entities, enhancing the ability to capture relationships, through innovations in constructing graphs. StratMed(Li et al., 2023) stratifies different disease/procedure-medication relationships, enhancing the representation of sparse relationships to alleviate the long-tail problem in the data.
The third category of methods involves incorporating partial medication-related knowledge to address safety and reliability issues in medication combinations. SafeDrug (Yang et al., 2021b) uses molecular data to improve medication relationship characterization and reduce adverse DDIs. MoleRec (Yang et al., 2023) simulates molecular substructures for better DDI representation. Carmen (Chen et al., 2023) integrates patient history with molecular learning to differentiate similar molecules by function.
@@ 2.2. Causal Inference in Recommender Systems
The incorporation of causal inference (Gao et al., 2024) into recommender systems(RS) still presents as a relatively novel concept, and ongoing research directions can be broadly categorized into three main categories. The primary objective of the first research category is to tackle data bias issues. For instance, methods such as the backdoor criterion employed by (Wei et al., 2021; Zheng et al., 2021a; Wang et al., 2022d) aim to eliminate confounding factors, addressing challenges like popularity bias and exposure bias in the dataset.
In the second category of work, exemplified by (Zhang et al., 2021; Xiao and Wang, 2022; Wang et al., 2022c), real-world data collection is coupled with inverse probability weighting and counterfactual methods. This involves introducing counterfactual interactions that contradict observations to fill data gaps or reduce noise.
Additionally, the third category of works focus on enhancing model interpretability and fairness. (Wang et al., 2022a; Si et al., 2022) introduce causal discovery and counterfactual approaches in RS to tackle the inherent opacity issues associated with deep learning models.
Our method utilizes causal inference to establish causal relationships with clear direction and precise quantification, effectively overcoming the inherent limitations of traditional methods that depend on co-occurrence relationships.
@@ 4.2. Relationship Mining
At this stage, we introduce methods of causal discovery and causal estimation to uncover the underlying genuine relationships among medical entities and eliminate spurious relationships caused by mere co-occurrence. We transform set-to-set relationships into precise point-to-point relationships and refine the causal effects for each group of relationships.
@@ 4.3. Health State Learning
At this stage, we utilize causal relationships to learn the health state of patients. In homomorphic relationships, we capture the meanings of entities under different health states to update their representations; In heterogeneous relationships, we apply point-to-point disease/procedure-medication causality to update entity representations. Ultimately, we integrate this information to create a personalized representation of the clinical visit.
@@ 4.3.1. Homomorphic Relationship Learning
As shown in the left half of the Health State Learning part of Figure 2. In this stage, we primarily learn the representations of various entities through the relationships between homomorphic entities. By exploring pathological relationships based on causality, we delve into the dynamic differences between entities to enhance our ability to capture the personalized characteristics of patients.
In different health states, the impact of the same disease on patients can vary. In one state, a disease may be the cause of other diseases, while in another, it may be the result of them. Therefore, we introduce a Dynamic Self-Adaptive Attention (DSA) mechanism to learn the role of diseases in various health states. By using causal positioning to determine the role of a disease in the current clinical setting and assigning different weights, we enhance the personalized representation of patients.
First, we extract relationships from the causal graph $G_{v_{t}}$ that exclusively involves entities from the current visit and construct three isomorphic causal graphs: $G^{d}_{v_{t}}$, $G^{p}_{v_{t}}$, and $G^{m}_{v_{t-1}}$. And then we categorize the patient’s disease/procedure, as shown in Figure 4. taking the example of diseases, for the learning of differences between diseases in a clinical visit, we categorize the diseases $\mathcal{D}_{t}$ into four types $\mathcal{D}^{j}_{t}$ based on $G^{d}_{v_{t}}$: 1) Causal diseases $\mathcal{D}^{1}_{t}$, which initiate other diseases and represent the primary ailment; 2) Effect diseases $\mathcal{D}^{2}_{t}$, influenced by other diseases and indicative of secondary symptoms; 3) Middle diseases $\mathcal{D}^{3}_{t}$, both influencing and influenced by other diseases; 4) Independent diseases $\mathcal{D}^{4}_{t}$, existing without direct causal links to other diseases.
Since each patient has a unique disease history and record, the same disease may occupy different positions in the causal graphs $G^{d}_{v_{t}}$ of different patients, leading to classification into different disease groups. The DSA mechanism learns the dynamic changes of disease groups in different health states, effectively capturing the personalized differences of diseases when facing different patients. It updates the representations of each entity based on these groupings to enhance the expressiveness of key entities. The specific formula is as follows:
@@
where $\text{Classify}(\cdot)$ assigns $d_{i}$ to a specific category $\mathcal{D}^{j}_{t}$, dynamic weight $w^{j}_{t}$ is generated by DSA, $\mathbf{W}$ and $b$ are trainable weight matrix and bias term, $\mathbf{h}^{D^{j}}_{t}$ represents the sum of embeddings of diseases within the ${D^{j}}_{t}$ set, and $\mathbf{h}^{r}_{d}$ is the modified embedding of diseases. We have extended the same approach to procedures and medications, obtained $\mathbf{h}^{r}_{p}$ and $\mathbf{h}^{r}_{m}$.
@@ 4.3.2. Heterogeneous Relationship Learning
where $\mathbf{W}_{r}^{l}$ is the weight matrix to relation $r$ at layer $l$, $\mathbf{W}_{0}^{l}$ is the weight matrix to node itself, $\mathcal{N}_{r}({d_{i}})$ is the set of neighboring nodes of ${d_{i}}$ with respect to relation $r$ and $c_{{d_{i}},r}$ is a normalization factor ensuring consistency in the influence of neighboring nodes under different relations. $I$ is the identity matrix, $\Theta_{r}^{l}$ is the weight update matrix for relation $r$ at layer $l$. In $G^{pm}_{v_{t}}$, we similarly repeat the aforementioned process, ultimately obtaining the entity embeddings $\mathbf{h}^{e}_{d}$, $\mathbf{h}^{e}_{p}$, and $\mathbf{h}^{e}_{m}$.
@@ 4.3.3. Aggregation of Entity Information
We obtain three set representations $\mathbf{h}_{\mathcal{D}_{t}}$, $\mathbf{h}_{\mathcal{P}_{t}}$, and $\mathbf{h}_{\mathcal{M}_{t-1}}$ by aggregating the updated embeddings with element-wise addition, which are then concatenated to the representation $\mathbf{h}_{v_{t}}$ for the current clinical visit.
@@ 4.4. Medication Recommendation
To deep mine personalized patient representations, we consider the patient’s medical history. We represent each medical record and integrate this information to recommend medication combinations. Using multiple GRUs to capture temporal dependencies among visits, we generate the GRU output $\mathbf{o}_{v_{t}}$ and patient embedding $\mathbf{h}_{\mathcal{H}}$. An MLP layer with activation function $\sigma$ converts this representation into medication scores. Medication $\hat{m}_{i}$ with a score exceeding $\delta$ will be chosen.

### v3

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USA.booktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAisbn: 979-8-4007-0436-9/24/10doi: 10.1145/3627673.3679542ccs: Information systems Information retrievalccs: Computing methodologies Causal reasoning and diagnosticsccs: Applied computing Health informatics
@@ 2.1. Medication Recommendation
Early works in medication recommendation, such as LEAP (Zhang et al., 2017), are instance-based, focusing on single-visit data and treating medication recommendation as a multi-instance, multi-label classification task.
The first category of research models sequential relationships in patients’ medical records to enhance personalization. RETAIN (Choi et al., 2016b) personalizes disease predictions using patient histories. MICRON (Yang et al., 2021a) optimizes therapy by adjusting medications for new symptoms. COGNet (Wu et al., 2022) translates diseases/procedures to medications and reuses effective past treatments.
The second category improves prescription accuracy by capturing relationships between entities. DMNC (Le et al., 2018) uses dynamic memory and deep learning to recommend precise medications. Studies like (Zheng et al., 2021b, 2023; Wu et al., 2023) use graph networks to describe connections between medical entities. StratMed (Li et al., 2023) stratifies relationships to alleviate data sparsity.
The third category addresses safety and reliability in medication combinations by incorporating partial medication-related knowledge. SafeDrug (Yang et al., 2021b) and MoleRec (Yang et al., 2023) use molecular data to reduce adverse DDIs. Carmen (Chen et al., 2023) integrates patient history with molecular learning to differentiate similar molecules by function.
@@ 2.2. Causal Inference in Recommender Systems
The incorporation of causal inference (Gao et al., 2024) into recommender systems (RS) is a novel concept, with ongoing research falling into three main categories.
The first category focuses on tackling data bias. Methods like the backdoor criterion (Wei et al., 2021; Zheng et al., 2021a; Wang et al., 2022d) aim to eliminate confounding factors, addressing challenges such as popularity and exposure bias.
The second category, exemplified by (Zhang et al., 2021; Xiao and Wang, 2022; Wang et al., 2022c), couples real-world data collection with inverse probability weighting and counterfactual methods to fill data gaps or reduce noise by introducing counterfactual interactions.
The third category aims to enhance model interpretability and fairness. Studies like (Wang et al., 2022a; Si et al., 2022) introduce causal discovery and counterfactual approaches in RS to address the opacity issues inherent in deep learning models.
Our method uses causal inference to establish clear, quantifiable causal relationships, effectively overcoming the limitations of traditional methods that rely on co-occurrence relationships.
@@ 4.2. Relationship Mining
At this stage, we use causal discovery and estimation to uncover genuine relationships among medical entities, eliminating spurious co-occurrence relationships. We transform set-to-set relationships into precise point-to-point relationships and refine causal effects for each group.
@@ 4.3. Health State Learning
We use causal relationships to learn patients’ health states. Homomorphic relationships update entity meanings under different health states, while heterogeneous relationships apply disease/procedure-medication causality. This information is integrated to create a personalized clinical visit representation.
@@ 4.3.1. Homomorphic Relationship Learning
As shown in the left half of the Health State Learning part of Figure 2, we learn entity representations through homomorphic relationships. By exploring pathological relationships based on causality, we capture the dynamic differences between entities, enhancing our ability to capture personalized patient characteristics.
In different health states, a disease’s impact on patients can vary. In one state, a disease may cause other diseases, while in another, it may result from them. We introduce a Dynamic Self-Adaptive Attention (DSA) mechanism to learn the role of diseases in various health states. Using causal positioning to determine a disease’s role in the current clinical setting, we assign different weights, enhancing personalized patient representations.
First, we extract relationships from the causal graph $G_{v_{t}}$ involving entities from the current visit and construct three isomorphic causal graphs: $G^{d}_{v_{t}}$, $G^{p}_{v_{t}}$, and $G^{m}_{v_{t-1}}$. We then categorize the patient’s diseases/procedures as shown in Figure 4. For diseases $\mathcal{D}_{t}$, we categorize them into four types $\mathcal{D}^{j}_{t}$ based on $G^{d}_{v_{t}}$: 1) Causal diseases $\mathcal{D}^{1}_{t}$, which initiate other diseases and represent the primary ailment; 2) Effect diseases $\mathcal{D}^{2}_{t}$, influenced by other diseases and indicative of secondary symptoms; 3) Middle diseases $\mathcal{D}^{3}_{t}$, both influencing and influenced by other diseases; 4) Independent diseases $\mathcal{D}^{4}_{t}$, existing without direct causal links to other diseases.
Since each patient has a unique disease history, the same disease may occupy different positions in the causal graphs $G^{d}_{v_{t}}$ for different patients, leading to different group classifications. The DSA mechanism learns the dynamic changes of disease groups across various health states, capturing personalized differences. It updates entity representations based on these groupings to enhance the expressiveness of key entities. The specific formula is as follows:
@@
Where $\text{Classify}(\cdot)$ assigns $d_{i}$ to a specific category $\mathcal{D}^{j}_{t}$, the DSA generates dynamic weight $w^{j}_{t}$. $\mathbf{W}$ and $b$ are trainable weight matrix and bias term, $\mathbf{h}^{D^{j}}_{t}$ represents the sum of embeddings of diseases within $\mathcal{D}^{j}_{t}$, and $\mathbf{h}^{r}_{d}$ is the modified embedding of diseases. The same approach is extended to procedures and medications, yielding $\mathbf{h}^{r}_{p}$ and $\mathbf{h}^{r}_{m}$.
@@ 4.3.2. Heterogeneous Relationship Learning
Where $\mathbf{W}_{r}^{l}$ is the weight matrix for relation $r$ at layer $l$, $\mathbf{W}_{0}^{l}$ is the weight matrix for the node itself, $\mathcal{N}_{r}({d_{i}})$ is the set of neighboring nodes of ${d_{i}}$ for relation $r$, and $c_{{d_{i}},r}$ is a normalization factor. $I$ is the identity matrix, and $\Theta_{r}^{l}$ is the weight update matrix for relation $r$ at layer $l$. In $G^{pm}_{v_{t}}$, we repeat this process to obtain entity embeddings $\mathbf{h}^{e}_{d}$, $\mathbf{h}^{e}_{p}$, and $\mathbf{h}^{e}_{m}$.
@@ 4.3.3. Aggregation of Entity Information
We use the residual method to stack the embeddings after initialization and weighting. By aggregating the updated embeddings with element-wise addition, we obtain three set representations: $\mathbf{h}_{\mathcal{D}_{t}}$, $\mathbf{h}_{\mathcal{P}_{t}}$, and $\mathbf{h}_{\mathcal{M}_{t-1}}$. These representations are then concatenated with the representation $\mathbf{h}_{v_{t}}$ for the current clinical visit.
@@ 4.4. Medication Recommendation
To deeply mine personalized patient representations, we consider the patient’s medical history. Each medical record is represented and integrated to recommend medication combinations. Multiple GRUs capture temporal dependencies among visits, generating GRU output $\mathbf{o}_{v_{t}}$ and patient embedding $\mathbf{h}_{\mathcal{H}}$. An MLP layer with activation function $\sigma$ converts this into medication scores. Medications $\hat{m}_{i}$ with scores exceeding $\delta$ are chosen.